OpenStudy (anonymous):

PLEASE HELP !! WILL GIVE MEDALS and become fan! cos 11pi/6 sin 5pi/4 tan 2pi/3 find the angle: -5pi/4 find the angle: -pi/3 sin -5pi/6 cos -7pi/4

OpenStudy (anonymous):

put the value of pi=180 and simplify

OpenStudy (anonymous):

multiple everything that has pi in it by 180? @duaali

OpenStudy (anonymous):

11*180/6 =330 and cos 330 = 0.866

OpenStudy (anonymous):

oh okay. I get it. can you help with some more?

OpenStudy (anonymous):

@duaali

OpenStudy (anonymous):

second one is sine 225 = -0.7071

OpenStudy (anonymous):

\[-1\div \sqrt{2}\]

OpenStudy (anonymous):

I understand those now, but need help with long division. I have six problems. (5x2+43x+56)/(x+7) (6m2-3m-30)/(2m-5) 36t2-39t+9)/(4t-3) (m2-7m-11)/(m-8) (n2-3n-21)/(n-7) (50x2+10x-35)/(5x-4) can you do step by step for these?

OpenStudy (anonymous):

u know the method of factoring so 5x^2+8x+35x+56=0 x(5x+8)+7(5x+8)=0 now u got two factors (5x+8)(x+7)

OpenStudy (anonymous):

(5x+8)(x+7)/(x+7) it'll give u 5X+8